A man applies a constant pulling Force on a heavy parcel of mass 50 kg using a light inextensible rope which process over a light frictionless pulley .The coefficient of static friction between the parcel in the rough table surface is 0.34 the magnitude of the maximum static friction force is 120 newtons. ignore the mass of the Rope
2.1 when the static friction force is at it's maximum show that the magnitude of the vertical component of the tension force in the Rope is 137.06
1
Expert's answer
2018-05-17T10:06:08-0400
Normal force is
N=F_fr/μ=120/0.34 N
The magnitude of the vertical component of the tension force in the Rope is
T_vertical=mg-N
T_vertical=50(9.8)-120/0.34=137.06 N.